Distance of a particle is given by x=40+12t-t³ then find the distance travel before coming to rest

Grade:11

1 Answers

Mahima Kanawat
6 years ago
In differentiating it with respect to time we will get the velocityV= 12~3t^2When it will come to rest V=0 On putting V = 03t^2= 12t = 2 sec On putting t = 2 Distance travelled will be X= 40 + 24 ~ 8= 56 m